# The Narrow Depth and Breadth of Corporate Responsible AI Research

### v1

@@ 3 Motivations for Industry to Engage in Responsible AI Research
AI firms, as the primary drivers of technological development and deployment, have a critical role to play in responsible AI research due to their unique position, expertise, and potential impact on society. Moreover, given that the design of AI necessitates considerable public deliberation, we contend that it is crucial for these organizations to publicly engage in responsible AI research.
@@
Unique position and capability. AI firms occupy a unique position to shape the trajectory of AI development due to their central role in developing and deploying this technology (Miric et al.,, 2023; Ahmed et al.,, 2023). The decisions these firms make during the design and deployment phases have far-reaching implications for society (Rudner,, 1953; Johnson,, 2023; Bijker and Law,, 1994). Furthermore, their privileged vantage point allows them to identify potential shortcomings and risks associated with their AI systems that may not be apparent to external researchers and take necessary measures to address those (Nedzhvetskaya and Tan,, 2022). This unique position imposes a heightened responsibility on AI firms to ensure that the technology they develop is socially beneficial and morally sound (Smith,, 2023).
@@
Access to required resources. The development of cutting-edge AI requires vast amounts of data, computational power, and talent–resources that are largely concentrated within AI firms (Ahmed et al.,, 2023; Birhane et al.,, 2023; Ahmed and Wahed,, 2020). Industry actors develop the vast majority of state-of-the-art AI models (Ahmed et al.,, 2023; Maslej et al.,, 2024), which are often opaque and difficult for outside researchers to reproduce and audit (Bommasani et al.,, 2023). Moreover, AI firms attract a disproportionate share of top AI talent, with nearly 70% of AI PhDs opting for industry roles over academic positions (Ahmed et al.,, 2023). Additionally, industry is dominating the leading AI benchmarks and directing the trajectory of AI research. This disparity in resources between industry and academia is particularly concerning given that certain limitations of AI models, such as toxicity in language, only become evident at larger scales (Ganguli et al.,, 2022; Birhane et al.,, 2023). This disparity underscores the heightened responsibility of AI firms to commercialize AI research responsibly.
@@
Cultivation of absorptive capacity. Engaging in responsible AI research is crucial for firms to develop their “absorptive capacity”–the ability to identify, assimilate, and apply external knowledge (Cohen and Levinthal,, 1990; Zahra and George,, 2002; Baruffaldi and Poege,, 2023). This process is complex and costly, requiring firms to establish specific routines and organizational procedures (Zahra and George,, 2002). Recent research suggests that active participation in conferences is essential for firms to effectively learn from other researchers (Baruffaldi and Poege,, 2023). By actively participating in responsible AI research and presenting their findings at leading conferences, AI firms can enhance their capacity to absorb and integrate the knowledge generated by the broader research community (Epstein et al.,, 2018; Microsoft,, 2024).
Increased transparency and accountability. Fostering transparency and accountability in the AI development process is crucial for developing trustworthy AI systems (Brundage et al.,, 2020; Mittelstadt,, 2019). Industry needs to engage in public deliberations and consult with external stakeholders when making critical decisions about design choices during AI development. This approach is essential for several reasons.
@@
First, participating in peer-review processes facilitates interaction with various stakeholders, particularly those in academia and the non-profit sectors. Such engagement is essential because mitigating bias and reducing unfairness in AI models often require making choices that conflict with each other (Raghavan,, 2023). Furthermore, the complexity of designing technology for diverse user needs, with varying incentives and often unclear user preferences, necessitates collaborative decision-making (Ghoshal and Dasgupta,, 2023; Kleinberg et al.,, 2023).
@@
Second, decisions regarding what to commercialize, how to commercialize, and for whom should undergo public deliberation (Barocas et al.,, 2020). It is equally important to consider what not to develop or commercialize as some AI technologies may have unintended negative consequences or raise significant ethical concerns (Cohen et al.,, 2024). Public deliberation is imperative because the repercussions of commercialization often disproportionately impact marginalized groups (Raji and Buolamwini,, 2019; Bruckner,, 2018). Transparency regarding the values encoded and reflected in AI models (Birhane et al.,, 2022) is crucial for developing AI systems that equitably represent and serve all individuals, as it enables stakeholders to identify and address potential biases or limitations that could lead to discriminatory or unfair outcomes. Furthermore, the substantial conceptual ambiguities and debates surrounding responsible AI (Powers and Ganascia,, 2020) require significant public deliberation with diverse stakeholders, including those from the Global South (Png,, 2022; Durmus et al.,, 2023). Stakeholders should be actively involved in determining the values that AI systems should align with and the risks and harms that need to be addressed (Diberardino et al.,, 2024). A sociotechnical approach emphasizes that no single group, especially technologists alone, should unilaterally make these critical decisions (Korinek and Balwit,, 2022). Making AI safe requires public debates on these questions (Lazar and Nelson,, 2023). By involving a wide range of perspectives and expertise, AI firms can develop more robust and ethically sound solutions to the complex problems they face (Neff,, 2020). Overall, increased responsible AI research by industry (and academia) will facilitate better communication within the AI research community, that prioritizes transparency and accountability, and leads to the development of trustworthy AI systems.
@@
Third, publishing in peer-reviewed forums facilitates greater scrutiny of the decision-making processes of AI firms. It also enables outsiders to observe how firms balance their business models and financial incentives with the direction of technology development (Zuboff,, 2015). This scrutiny is essential to ensure transparency and accountability in AI development (Brundage et al.,, 2020; Mittelstadt,, 2019) as it helps to identify potential conflicts of interest and encourages responsible innovation.
@@
Risk reduction. AI is a tool that lacks autonomy in decision-making. Therefore, just as automobile manufacturers, who are held liable for accidents due to vehicle defects and therefore undertake thorough pre-market testing, AI companies should also bear responsibility for the consequences of their products (Martin,, 2019). Engaging in rigorous responsible AI research is crucial for these companies to understand the limitations of their technology and products (Holstein et al.,, 2019).
@@
Financial benefits. Investing in responsible AI research yields substantial financial benefits for firms. By engaging in responsible AI research, industry can improve the quality of AI products and services by incorporating ethical considerations and ensuring the development of AI systems that are fair, unbiased, and trustworthy. Moreover, by developing internal expertise in this area, firms will be better equipped to navigate the complex challenges that will inevitably arise as the technology advances (Renieris et al.,, 2022). Collaborating with the broader responsible AI community and sharing knowledge will help industry leaders stay attuned to evolving societal concerns and adapt accordingly. Overall, firms that prioritize responsible AI will be more resilient and successful in the long run, as regulators and customers are more likely to trust products that are not prone to hallucinations and biases.
@@ 4 The Narrow Depth of Industry’s Responsible AI Research
4.1 Assessing Responsible AI Research Engagement of Industry and Academia
In this section, we provide a comprehensive analysis of the engagement of industry and academia in responsible AI research. We begin by evaluating the extent to which industry participates in responsible AI research and compare this engagement to their involvement in conventional AI research. This comparison allows us to understand the relative importance placed on responsible AI within industry’s research portfolio. Subsequently, we contrast industry’s commitment to responsible AI with academia’s research efforts in this domain. We consider both the publication count and the quality of the publications to gain a thorough understanding of the current state of responsible AI research in these two sectors.
@@ 4.1.1 Data Description
To identify AI firms and assess industry engagement in responsible AI research, we focused on companies holding at least one AI patent registered with the United States Patent and Trademark Office (USPTO), yielding a list of 1,771 AI firms. Patents signal a firm’s ability to actively research and commercialize technology in a field (Arora et al.,, 2018). Our sample included firms with patents under the Cooperative Patent Classification (CPC) class “computer systems based on specific computational models,” excluding quantum computing, per consultations with USPTO examiners. Afterward, we collected each AI firm’s peer-reviewed publications from Scopus, including publications from both conferences and journals. Our sampling strategy ensures that we have the full research portfolio of AI firms. This process resulted in 679,919 papers from 2010 to 2022, which included AI papers, non-AI papers, and responsible AI papers (dataset 1).
To compare industry research engagement with academia, we first collected data from Scopus for the leading 100 universities in computer science research. We selected these universities based on their research productivity between 2010 and 2022, as reported by CSRankings.org [^3] CSRankings (csrankings.org) is a well-cited resource in the literature (Ahmed et al.,, 2023; Meho,, 2019; Gofman and Jin,, 2023) that ranks universities based on their research output in computer science.. The complete list of universities can be found in Table LABEL:table:oldA2 in the appendix. For each of the selected universities, we downloaded their publication data from Scopus, covering the years 2010 to 2022. This process resulted in a comprehensive dataset (dataset 1) consisting of 5,265,419 peer-reviewed publications. This dataset serves as the basis for our analysis of academic research engagement.
@@ 4.1.2 Classification Method
The classification of responsible AI papers presents a significant challenge due to the lack of a universally accepted definition of responsible AI (Jobin et al.,, 2019). Consequently, researchers may employ varying definitions and classification approaches, leading to inconsistencies in the identification and categorization of responsible AI research. For our analysis, we conducted a supervised machine learning method using an ensemble classification (Dietterich,, 2000) that combined a transformer-based pre-trained SciBERT model (Beltagy et al.,, 2019) and a boosting-based XGBoost classifier (Chen and Guestrin,, 2016). Such ensembled classifiers are generally used to gain better results than those from the constituent classifiers alone (Caruana et al.,, 2004; Tulyakov et al.,, 2008). The intuition here is that our SciBERT model, combined with a separate classifier and being trained on research articles, would be able to learn the patterns from the data and identify responsible AI and conventional AI papers (see Fig. S1 in the appendix for our overall workflow).
@@
Transformer-based models outperform other neural network models on text processing tasks (Vaswani et al.,, 2017). Among them, BERT (Devlin et al.,, 2018) and other BERT-based models, like RoBERTa (Liu et al.,, 2019), have demonstrated state-of-the-art performance on different NLP tasks. For our analysis, we evaluated multiple pre-trained transformer-based models on a hold-out test sample. Our findings show the SciBERT model, originally trained on over a million research articles, outperformed other models, yielding better accuracy (see Table 2). [^4] As part of the ensemble model, we fine-tuned the SciBERT model by setting 50 epochs with 10-fold cross-validation. Additionally, we enabled early stopping with a delta of $1e^{-4}$, a patience value of 3 to avoid overfitting, and set the training batch size to 50.
@@
To complement the SciBERT model, we used the XGBoost classifier as it has demonstrated better performance on diverse data types when ensembled with a neural network-based model (Shwartz-Ziv and Armon,, 2022; Li et al.,, 2017).
Fine-tuning the ensemble model also included manual validation of the training dataset and feature ablation on it. Ablation studies are helpful in identifying and quantifying the actual impact of the constituent model components (Sheikholeslami et al.,, 2021). Following prior literature, we performed ablations on features (Merrick,, 2019; Girshick et al.,, 2014; Richardson et al.,, 2006) and training data subsets (Mousavi et al.,, 2020; Sheikholeslami et al.,, 2021). We tested multiple models by using varied subsets of features (e.g., abstract, title and abstract, abstract and publication outlet title). This careful ablation led us to a robust model which included paper title, paper abstract, and publication outlet title as the most effective features.
@@ 4.1.4 Findings
In Fig. 1a, we visualize firms based on their responsible AI publication engagement. First, we counted the number of AI firms (n = 1,771) based on their level of engagement in responsible AI research within the years 2010-22 and disaggregated them into three groups: (a) firms with no engagement in responsible AI research, (b) firms with limited engagement in responsible AI research (publishing fewer than five papers), and (c) firms with engagement in responsible AI research (publishing five or more papers).
@@
Figure 1: Here, Fig. 1a shows the number of firms that have conventional and responsible AI research engagement, divided into three groups based on their level of engagement. Fig. 1b showcases a stark difference in the leading 100 AI firms’ (n = 506,017 papers, 2010-22) and universities’ (n = 5,265,419 papers, 2010-22) participation in conventional AI research compared to their engagement in responsible AI research. Fig. 1c presents a citation-weighted paper count for the same data. A trend line reflects the participation trend within each group with 95% confidence intervals. In both Fig. 1b and 1c, the dashed lines indicate a reference line where the proportion of responsible AI papers to conventional AI papers is 2.5%.
Fig. 1a shows that an overwhelming majority of AI firms (89.9%) did not engage in any responsible AI research from 2010 to 2022. Additionally, only 3.3% of the AI firms had five or more responsible AI papers, indicating an unenthusiastic participation. In other words, 96.7% of AI firms published fewer than five responsible AI papers over 13 years.
This pattern of limited industry engagement continues for AI research firms, i.e., firms that had at least one AI research paper. We disaggregated papers from 519 such firms in the same groups as before and observed a majority of them (65.5%) did not engage in responsible AI research. Almost a fourth of them (23.3%) published fewer than five papers over 13 years, showing very limited participation. A mere 11.2% of these firms demonstrated meaningful engagement in responsible AI research. Overall, industry engagement in responsible AI research was limited among both AI patent-holding firms and AI research firms.
@@
Here, we compared the 100 leading academic research institutes with the 100 leading AI firms (for a comparison with our full sample of AI firms, see Fig. S2 in the appendix). We selected the leading AI research institutes and firms based on their number of conventional AI papers. This selection ensured a fair comparison in terms of research capability and reputation. Notably, the leading 100 AI firms have significantly greater resources than even the most elite universities (Ahmed et al.,, 2023). Fig. 1b indicates that industry’s participation in conventional AI research is on par with that of leading academic institutions. However, the leading AI firms engage in responsible AI research at a far lower rate than comparable leading academic institutions. This disparity is particularly visible from the slope of industry’s regression line, which is flatter than that of academia.
Finally, to take into account the quality of the publications, we used a citation-weighted publication count (see Section 4.1.3 for the details). Fig 1c suggests that after considering the quality of the publications, the disparity between industry and academia stayed largely similar. This contrast is visible from the slope of industry’s regression line, compared with academia’s regression line. Overall, this analysis shows that industry’s engagement in responsible AI research is limited even when we consider the quality of research. Interestingly, the performance of some AI firms has improved when quality was taken into account, suggesting that these firms produce high-quality responsible AI research, albeit in smaller quantities compared to their conventional AI research output.
Consistent with the findings of de Laat (2021), we observed that a smaller number of AI firms are indeed engaging in responsible AI research. However, the extent of their engagement is still significantly lower than their conventional AI research. When considering the quality of their contributions, we note that a smaller number of firms surpassed the 2.5% threshold. It is plausible that firms selectively disclose their high-quality papers, which may partially explain the observed improvement in performance when quality is accounted for. Nevertheless, the overall disparity between industry and academia in terms of responsible AI research output remains noticeable. This stark contrast suggests that market incentives to enhance AI capabilities far outweigh the incentives to prioritize and invest in responsible AI development.

### v2

@@ 3 Motivations for Industry to Engage in Responsible AI Research
AI firms, as primary drivers of technological development and deployment, hold a distinctive position in responsible AI research given their expertise, resources, and potential societal impact. Several theoretical perspectives suggest reasons why these organizations may benefit from publicly engaging in responsible AI research. We outline these perspectives below, recognizing that the actual motivations and engagement patterns of firms remain empirical questions.
@@
Unique position and capability. AI firms occupy a unique position to shape the trajectory of AI development due to their central role in developing and deploying this technology (Miric, Jia, and Huang, 2023; Ahmed, Wahed, and Thompson, 2023). The decisions these firms make during the design and deployment phases have far-reaching implications for society (Rudner, 1953; Johnson, 2023; Bijker and Law, 1994). Furthermore, their privileged vantage point allows them to identify potential shortcomings and risks associated with their AI systems that may not be apparent to external researchers and take necessary measures to address those (Nedzhvetskaya and Tan, 2022). This unique position imposes a heightened responsibility on AI firms to ensure that the technology they develop is socially beneficial and morally sound (Smith, 2023).
@@
Access to required resources. The development of cutting-edge AI requires vast amounts of data, computational power, and talent, resources that are largely concentrated within AI firms (Ahmed, Wahed, and Thompson, 2023; Birhane, Prabhu, Han, and Boddeti, 2023; Ahmed and Wahed, 2020). Industry actors develop the majority of state-of-the-art AI models (Ahmed, Wahed, and Thompson, 2023; Maslej, Fattorini, Perrault, Parli, Reuel, Brynjolfsson, Etchemendy, Ligett, Lyons, Manyika, Niebles, Shoham, Wald, and Clark, 2024), which are often difficult for external researchers to reproduce and audit (Bommasani, Klyman, Longpre, Kapoor, Maslej, Xiong, Zhang, and Liang, 2023). Moreover, AI firms attract a substantial share of top AI talent, with nearly 70% of AI PhDs opting for industry roles over academic positions (Ahmed, Wahed, and Thompson, 2023). Research suggests that certain limitations of AI models, such as toxicity in language, may only become evident at larger scales (Ganguli, Hernandez, Lovitt, Askell, Bai, Chen, Conerly, Dassarma, Drain, Elhage, El Showk, Fort, Hatfield-Dodds, Henighan, Johnston, Jones, Joseph, Kernian, Kravec, Mann, Nanda, Ndousse, Olsson, Amodei, Brown, Kaplan, McCandlish, Olah, Amodei, and Clark, 2022; Birhane, Prabhu, Han, and Boddeti, 2023). This concentration of resources suggests that industry engagement in responsible AI research may offer access to settings and scale that academic researchers face challenges replicating independently.
@@
Cultivation of absorptive capacity. Engaging in responsible AI research is crucial for firms to develop their “absorptive capacity”–the ability to identify, assimilate, and apply external knowledge (Cohen and Levinthal, 1990; Zahra and George, 2002; Baruffaldi and Poege, 2023). This process is complex and costly, requiring firms to establish specific routines and organizational procedures (Zahra and George, 2002). Recent research suggests that active participation in conferences is essential for firms to effectively learn from other researchers (Baruffaldi and Poege, 2023). By actively participating in responsible AI research and presenting their findings at leading conferences, AI firms can enhance their capacity to absorb and integrate the knowledge generated by the broader research community (Epstein, Payne, Shen, Dubey, Felbo, Groh, Obradovich, Cebrian, and Rahwan, 2018; Microsoft, 2024).
Transparency and accountability. Prior literature suggests that transparency and accountability in AI development may contribute to the trustworthiness of AI systems (Brundage, Avin, Wang, Belfield, Krueger, Hadfield, Khlaaf, Yang, Toner, Fong, Maharaj, Koh, Hooker, Leung, Trask, Bluemke, Lebensold, O’Keefe, Koren, and Ryffel, 2020; Martin, 2019b; Mittelstadt, 2019). Martin (2019b) argues that algorithms are value-laden rather than neutral, creating moral consequences and influencing stakeholder rights, which grounds firm accountability regardless of algorithmic complexity. Engagement in public deliberations and consultation with external stakeholders represents one mechanism through which firms may navigate critical decisions about design choices during AI development. Several dimensions of this relationship have been explored in the literature.
@@
First, participating in peer-review processes facilitates interaction with various stakeholders, particularly those in academia and the non-profit sectors. Research indicates that mitigating bias and reducing unfairness in AI models often require making choices that conflict with each other (Raghavan, 2023). Furthermore, the complexity of designing technology for diverse user needs, with varying incentives and often unclear user preferences, may benefit from collaborative decision-making (Ghoshal and Dasgupta, 2023; Kleinberg, Mullainathan, and Raghavan, 2023).
@@
Second, decisions regarding what to commercialize, how to commercialize, and for whom have been identified as candidates for public deliberation (Barocas, Biega, Fish, Niklas, and Stark, 2020). Scholars have also emphasized what not to develop or commercialize, as some AI technologies may have unintended negative consequences or raise significant ethical concerns (Cohen, Kolt, Bengio, Hadfield, and Russell, 2024). This perspective reflects concerns that the repercussions of commercialization may disproportionately impact marginalized groups (Raji and Buolamwini, 2019; Bruckner, 2018). Transparency regarding the values encoded and reflected in AI models (Birhane, Kalluri, Card, Agnew, Dotan, and Bao, 2022) may enable stakeholders to identify potential biases or limitations in AI systems. Furthermore, the substantial conceptual ambiguities and debates surrounding responsible AI (Powers and Ganascia, 2020) highlight the potential value of public deliberation with diverse stakeholders, including those from the Global South (Png, 2022; Durmus, Nyugen, Liao, Schiefer, Askell, Bakhtin, Chen, Hatfield-Dodds, Hernandez, Joseph, Lovitt, McCandlish, Sikder, Tamkin, Thamkul, Kaplan, Clark, and Ganguli, 2023). Recent scholarship emphasizes the role of diverse stakeholders in determining the values that AI systems align with and the risks and harms to be addressed (Diberardino, Baleshta, and Stark, 2024). A sociotechnical approach emphasizes that no single group, especially technologists alone, is positioned to unilaterally make these critical decisions (Korinek and Balwit, 2022). Some scholars argue that public debates on these questions contribute to AI safety (Lazar and Nelson, 2023). Involving a range of perspectives and expertise may support the development of more robust solutions to complex challenges (Neff, 2020). Overall, research engagement in responsible AI may facilitate communication within the AI research community around transparency and accountability.
@@
Third, publishing in peer-reviewed forums facilitates greater scrutiny of the decision-making processes of AI firms. It also enables outsiders to observe how firms balance their business models and financial incentives with the direction of technology development (Zuboff, 2015). Such scrutiny may support transparency and accountability in AI development (Brundage, Avin, Wang, Belfield, Krueger, Hadfield, Khlaaf, Yang, Toner, Fong, Maharaj, Koh, Hooker, Leung, Trask, Bluemke, Lebensold, O’Keefe, Koren, and Ryffel, 2020; Mittelstadt, 2019).
@@
Risk reduction. AI is a tool that lacks autonomy in decision-making. Drawing an analogy to automobile manufacturers, who are held liable for accidents due to vehicle defects and therefore undertake thorough pre-market testing, some scholars have argued that AI companies may similarly bear responsibility for the consequences of their products (Martin, 2019a). From this perspective, engagement in responsible AI research may help firms better understand the limitations of their technology and products (Holstein, Wortman Vaughan, Daumé, Dudik, and Wallach, 2019).
@@
Potential benefits. Literature suggests that investment in responsible AI research may yield benefits for firms. By engaging in responsible AI research, firms may improve the quality of AI products and services by incorporating ethical considerations and developing AI systems that are fair, unbiased, and trustworthy. Moreover, by developing internal expertise in this area, firms may be better positioned to navigate the complex challenges that arise as the technology advances (Renieris, Kiron, and Mills, 2022). Collaborating with the broader responsible AI community and sharing knowledge may help industry leaders stay attuned to evolving societal concerns and adapt accordingly. Research in this area suggests that firms with demonstrated commitment to responsible AI may benefit from increased trust among regulators and customers (Renieris, Kiron, and Mills, 2022). Additionally, firms that demonstrate commitment to responsible AI principles may be better positioned to attract and retain top AI talent, who increasingly prioritize ethical considerations when selecting employers (Ahmed and Jia, 2023).
@@ 4 The Narrow Depth of Industry’s Responsible AI Research
4.1 Comparing Responsible AI Research Output Across Industry and Academia
In this section, we provide a comprehensive analysis of responsible AI research output from industry and academia. We begin by examining the extent to which industry publishes responsible AI research and compare this output to their involvement in conventional AI research. This comparison allows us to characterize the composition of responsible AI research within industry’s research portfolio. Subsequently, we compare industry’s responsible AI research output with academia’s research output in this domain. We consider both publication counts and publication quality to document the current distribution of responsible AI research across these two sectors.
@@ 4.1.1 Data Description
To identify AI firms and assess industry engagement in responsible AI research, we focused on companies holding at least one AI patent registered with the United States Patent and Trademark Office (USPTO), yielding a list of 1,771 AI firms. Patents signal a firm’s ability to actively research and commercialize technology in a field (Arora, Belenzon, and Patacconi, 2018). Our sample included firms with patents under the Cooperative Patent Classification (CPC) class “computer systems based on specific computational models,” excluding quantum computing, per consultations with USPTO examiners. Afterward, we collected each AI firm’s peer-reviewed publications from Scopus, including publications from both conferences and journals. Our sampling strategy ensures that we have the full research portfolio of AI firms. This process resulted in 679,919 papers from 2010 to 2022, which included AI papers, non-AI papers, and responsible AI papers (dataset 1).
To compare industry research engagement with academia, we first collected data from Scopus for the leading 100 universities in computer science research. We selected these universities based on their research productivity between 2010 and 2022, as reported by CSRankings.org [^3] CSRankings (csrankings.org) is a well-cited resource in the literature (Ahmed, Wahed, and Thompson, 2023; Meho, 2019; Gofman and Jin, 2023) that ranks universities based on their research output in computer science.. The complete list of universities can be found in Table LABEL:table:oldA2 in the appendix. For each of the selected universities, we downloaded their publication data from Scopus, covering the years 2010 to 2022. This process resulted in a comprehensive dataset (dataset 1) consisting of 5,265,419 peer-reviewed publications. This dataset serves as the basis for our analysis of academic research engagement.
@@ 4.1.2 Classification Method
The classification of responsible AI papers presents a significant challenge due to the lack of a universally accepted definition of responsible AI (Jobin, Ienca, and Vayena, 2019). Consequently, researchers may employ varying definitions and classification approaches, leading to inconsistencies in the identification and categorization of responsible AI research. For our analysis, we conducted a supervised machine learning method using an ensemble classification (Dietterich, 2000) that combined a transformer-based pre-trained SciBERT model (Beltagy, Lo, and Cohan, 2019) and a boosting-based XGBoost classifier (Chen and Guestrin, 2016). Such ensembled classifiers are generally used to gain better results than those from the constituent classifiers alone (Caruana, Niculescu-Mizil, Crew, and Ksikes, 2004; Tulyakov, Jaeger, Govindaraju, and Doermann, 2008). The intuition here is that our SciBERT model, combined with a separate classifier and being trained on research articles, would be able to learn the patterns from the data and identify responsible AI and conventional AI papers (see Fig. S1 in the appendix for our overall workflow).
@@
Transformer-based models outperform other neural network models on text processing tasks (Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin, 2017). Among them, BERT (Devlin, Chang, Lee, and Toutanova, 2018) and other BERT-based models, like RoBERTa (Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis, Zettlemoyer, and Stoyanov, 2019), have demonstrated state-of-the-art performance on different NLP tasks. For our analysis, we evaluated multiple pre-trained transformer-based models on a hold-out test sample. Our findings show the SciBERT model, originally trained on over a million research articles, outperformed other models, yielding better accuracy (see Table 2). [^4] As part of the ensemble model, we fine-tuned the SciBERT model by setting 50 epochs with 10-fold cross-validation. Additionally, we enabled early stopping with a delta of $1e^{-4}$, a patience value of 3 to avoid overfitting, and set the training batch size to 50.
@@
To complement the SciBERT model, we used the XGBoost classifier as it has demonstrated better performance on diverse data types when ensembled with a neural network-based model (Shwartz-Ziv and Armon, 2022; Li, Situ, Gao, Yang, and Liu, 2017).
Fine-tuning the ensemble model also included manual validation of the training dataset and feature ablation on it. Ablation studies are helpful in identifying and quantifying the actual impact of the constituent model components (Sheikholeslami, Meister, Wang, Payberah, Vlassov, and Dowling, 2021). Following prior literature, we performed ablations on features (Merrick, 2019; Girshick, Donahue, Darrell, and Malik, 2014; Richardson, Prakash, and Brill, 2006) and training data subsets (Mousavi, Khanal, and Estrada, 2020; Sheikholeslami, Meister, Wang, Payberah, Vlassov, and Dowling, 2021). We tested multiple models by using varied subsets of features (e.g., abstract, title and abstract, abstract and publication outlet title). This careful ablation led us to a robust model which included paper title, paper abstract, and publication outlet title as the most effective features.
@@ 4.1.4 Findings
Figure 1: Here, Fig. 1a showcases the distribution of the leading 100 AI firms’ (n = 506,017 papers, 2010-22) and universities’ (n = 5,265,419 papers, 2010-22) participation in conventional AI research compared to their engagement in responsible AI research. Fig. 1b presents a citation-weighted paper count for the same data. A trend line reflects the participation trend within each group with 95% confidence intervals. In both Fig. 1a and 1b, the dashed lines indicate a reference line where the proportion of responsible AI papers to conventional AI papers is 2.5%.
We begin by examining patterns of responsible AI research engagement among AI research firms (those with at least one AI publication). Among 519 such firms, we observe considerable heterogeneity in publication activity: 34.5% engaged in responsible AI research during the study period (2010–2022), while 65.5% had no responsible AI publications. Among firms that did engage, output varied substantially, with a subset demonstrating sustained commitment through five or more publications. This variation may reflect differences in firm capabilities, strategic priorities, or organizational factors that shape research investment decisions. These patterns provide a baseline for examining how industry engagement in responsible AI compares to conventional AI research output.
@@
Here, we compared the 100 leading academic research institutes with the 100 leading AI firms (for a comparison with our full sample of AI firms, see Fig. S2 in the appendix). We selected the leading AI research institutes and firms based on their number of conventional AI papers. This selection ensured a fair comparison in terms of research capability and reputation. Notably, the leading 100 AI firms have significantly greater resources than even the most elite universities (Ahmed, Wahed, and Thompson, 2023). Fig. 1a indicates that industry’s participation in conventional AI research is comparable to that of leading academic institutions. However, on average, leading AI firms exhibit lower rates of responsible AI research output than comparable leading academic institutions. This difference is reflected in the slope of industry’s regression line, which is flatter than that of academia. Notably, we observe considerable heterogeneity across firms: several organizations, including IBM, Google, and Microsoft, demonstrate responsible AI publication rates comparable to elite research universities, and the number of firms with active responsible AI research programs has increased since 2018. These patterns suggest variation in how different firms allocate research effort between conventional and responsible AI topics.
Finally, to account for the quality of publications, we used a citation-weighted publication count (see Section 4.1.3 for details). Fig 1b suggests that after considering publication quality, the difference between industry and academia remains similar in magnitude. This pattern is observable in the slope of industry’s regression line compared with academia’s regression line. Notably, all papers in our analysis are peer-reviewed, ensuring comparable quality standards across sectors. The performance of some AI firms improved when quality was taken into account, suggesting that these firms produce high-impact responsible AI research, albeit in smaller quantities relative to their conventional AI output. Indeed, several industry-authored responsible AI papers rank among the most highly cited in the field, and academic researchers increasingly cite industry responsible AI work (see Fig. S20), indicating that the research quality of active industry contributors is recognized within the broader scholarly community.
Consistent with the findings of de Laat (2021), we observed that a subset of AI firms publish responsible AI research. However, their responsible AI output remains lower in proportion to their conventional AI research. When considering the quality of their contributions, we note that a smaller number of firms surpassed the 2.5% threshold. It is plausible that firms selectively disclose their high-quality papers, which may partially explain the observed improvement in performance when quality is accounted for. Overall, the difference between industry and academia in terms of responsible AI research output is persistent across our measures. This pattern suggests that variation exists in how market incentives shape research portfolio composition across organizational types (Ahmed et al., 2026).